find the distance between (-2,9) and (-5,15)
a) 18.67
b) 45.02
c) 6.71
d) 24.18

Answer by checkley77(12844)

(-2,9) and (-5,15)
X d9stance=-2+5=3 units.
Y distance=15-9=6 units.
a^2+b^2=c^2
3^2+6^2=c^2
9+36=c^2
c^2=45
c=sqrt45
c=6.71 ans. C) for the distance.